40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1353     Entry time: Wed Mar 4 03:50:17 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Still won't go above arm power 10 
Yoichi, Kentaro

The IFO still loses lock at around arm power 10.
I attached time series of various error signals when losing lock.
In all of the three cases, both of the arm powers go up rapidly just before losing lock.
(The first attachment was taken before I fixed the QPDX switching, so you can see saturation in TRX.)
But PD1_DC (the DC power in the PRC) did not go up in the third case.

I should also check correlations with laser power, CARM length (OSEM signals), seismic noise etc.
Attachment 1: lockLoss1.png  18 kB  | Hide | Hide all
lockLoss1.png
Attachment 2: lockLoss2.png  19 kB  | Hide | Hide all
lockLoss2.png
Attachment 3: lockLoss3.png  20 kB  | Hide | Hide all
lockLoss3.png
ELOG V3.1.3-